OpenStudy (anonymous):

(x-1)^2+(y-2)^2=9 how do I find the graphing points for this circle?

OpenStudy (anonymous):

The equation of a circle is: \[(x-h)^2+(y-k)^2=r^2\](h,k) is the center of the circle. Can you find the center?

OpenStudy (anonymous):

Thank you Trainwreaking, but could you go into more detail about how to work the problem out? Mathematics is not my strong suit...

OpenStudy (anonymous):

and the center would be -1,-2 right?

OpenStudy (anonymous):

Not quite. Because the original equation of the circle already has negative signs at (x-h) and (y-k), the 'h' and the 'k' don't get the negatives. x-h=x-1 and y-k=y-2 Here, you need to solve for the h and the k. x-h=x-1 -h=-1 h=1 y-k=y-2 -k=-2 k=2 So the center would be at (1,2).

OpenStudy (anonymous):

Now, you need to find the other points so that you can actually graph the circle. In the given equation, you are given 9, which is the r^2, the radius squared, in the equation of a circle. If \[9=r^2\] can you find the radius?

OpenStudy (anonymous):

R=3^2 right?

OpenStudy (anonymous):

\[r^2=9\] So you need to take the square root of both sides. \[\sqrt{r^2}=\sqrt{9}\]r=√9 r=3

OpenStudy (anonymous):

okay, I understand that.